OpenStudy (anonymous):

By how much does the mean of the data set below change if the outlier is removed? Round your answer to the nearest tenth, if necessary. 8, 6, 9, 12, 19, 7, 10, 8 (1 point) 0.5 0.6 1.3 2.4

OpenStudy (anonymous):

@phi @satellite73 @amistre64 @saifoo.khan @kropot72

OpenStudy (anonymous):

@sami-21 @electrokid

OpenStudy (anonymous):

HELP ME SOMEBODY PLEASE

OpenStudy (zehanz):

First, calulate the mean: m1= (8 + 6 + 9 + 12 + 19 + 7 + 10 + 8)/8. Then remove the outlier (19) Now calculate the mean again: m2= (8 + 6 + 9 + 12 + 7 + 10 + 8)/7. If you then calculate m1-m2, you've found the difference.

OpenStudy (anonymous):

is it 0.6?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

No, I've got another result. What results did you get for both means?

OpenStudy (anonymous):

72 and 372

OpenStudy (zehanz):

Both are wrong. m1 = (8 + 6 + 9 + 12 + 19 + 7 + 10 + 8)/8. So: add up all the number, then divide by 8. You may use your calculator!

OpenStudy (anonymous):

79/8 @ZeHanz

OpenStudy (zehanz):

=9.875 OK, just write it as a decimal number. The answer options are also.

OpenStudy (zehanz):

What is your outcome of m2 = (8 + 6 + 9 + 12 + 7 + 10 + 8)/7 ?

OpenStudy (anonymous):

60/7 @ZeHanz

OpenStudy (zehanz):

yes, that is 8.5714... Now the difference is 9.875-8.5714

OpenStudy (anonymous):

1.3 !!!! thanks for all your help

OpenStudy (zehanz):

YW!